sin(x)=1/2; -pi/2<x<pi/2; find the x
1
Expert's answer
2020-08-10T17:43:21-0400

sin(x) = 1/2

sin(x) = sin("\pi"/6)

x = "\pi"/6

"\pi"/6 { -"\pi"/2<x<"\pi"/2 }